(1 point) Book Problem 24 * F(z) = *5(e)dt , where f(t) = 51 9+u? du, find u F(x) = and F(3)
Answer #1

Y JF Fon) - S. f(t) dt where, f(t) = J9 +4² du 13 ų F() 1. FCt) dit 13 By using newton leibtnir sule I til forces de) h (7cm)F(n) = 9 +(025) 9 +1% d (1) d (25) dr NS 1 da 10 Fcom) 9+ or A (5044) N10 (0) no 1 F”(0) 59+222 Answer r F) 59+ 310 3 5.13